Given that 1, x and 3 are consecutive terms of an A.P. and 1, y and x are consecutive terms of a G.P., find the possible values of x and y.
Answer by richwmiller(17219)

3-x=x-1
4=2x
2=x
x/y=y/1
x=y^2
y=+-sqrt(2)
1,y,x
1,+-sqrt(2),2